40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 20 20:20:36 2020, gautam, Update, PSL, PMC servo checkout PMCsweep.pdfOLTFmeas.pdf
    Reply  Wed Jan 22 11:22:39 2020, gautam, Update, PSL, PMC modulation depth measurement modDepth.pdf
       Reply  Wed Jan 22 20:12:36 2020, gautam, Update, PSL, PMC demodulator electrical characterization demodChar.pdfmixerChar.pdf
          Reply  Thu Jan 23 14:37:05 2020, gautam, Update, PSL, PMC VGA chip damaged? VGAchar.pdf
             Reply  Thu Jan 23 16:37:14 2020, rana, Update, PSL, PMC VGA chip damaged? 
             Reply  Thu Jan 23 22:10:01 2020, gautam, Update, PSL, PMC servo pulled out 
                Reply  Fri Jan 24 15:42:08 2020, gautam, Update, PSL, PMC servo restored 
                   Reply  Sun Jan 26 13:47:00 2020, gautam, Update, PSL, PMC servo characterization elecTFs.pdfVGAchar_postFix.pdfVGAlinearity_postFix.pdfnewOLTFs.pdf
                      Reply  Tue Jan 28 14:33:24 2020, gautam, Update, PSL, Inferred free-running frequency noise inLoopNoise_IMClocked.pdffreqNoiseComparison.pdf
          Reply  Thu Jan 23 18:52:31 2020, gautam, Update, PSL, PMC RFPD characterization PDresp.pdf
Message ID: 15144     Entry time: Thu Jan 23 14:37:05 2020     In reply to: 15143     Reply to this: 15146     15149
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC VGA chip damaged? 

[jordan, gautam]

Summary:

The AD602 chip which implements the overall servo gain for the PMC seems to be damaged. We should switch this out at the next opportunity.

Details:

  1. According to the PSL cross connect wiring diagram, the VME DAC that provides the control voltage to the VGA stage goes to pins 7/8 of cross connect J16.
    • Jordan and I verified that the voltage at this point [Vout], is related to the PMC_GAIN EPICS slider [dB] value according to the following relationship: V_{\mathrm{out}} = (10-\mathrm{dB})/2.
  2. On the PMC servo board, this voltage is scaled by a factor of -1/10. 
    • This was confirmed by peeking at this voltage using a DMM (I clipped onto R31) while the gain slider was varied.
    • This corresponds to +/- 1000 mV reaching the AD602.
    • However, the AD602 is rated to work with a control voltage varying between +/- 625 mV.
    • What this means is that the EPICS slider value is not the gain of the AD602 stage. The latter is given by the relation G [\mathrm{dB}] = 32 \times V_{\mathrm{G}} + 10.
    • @team PSL upgrade: this should be fixed in the database file for the new c1psl machine.
  3. Using TP1 and TP2 connected to the SR785, I measured the transfer function of the AD602 for various values of the EPICS slider.
    • Result is shown in Attachment #1.
    • I did this measurement with the PMC locked, so I'm using the in-loop error signal to infer the gain of the VGA stage.
    • As expected, the absolute value of the gain does not match that of the EPICS slider (note that the AD602 has an input impedance of 100 ohms. So the 499 ohm series resistor between TP1 and the input of AD602 makes a 1/5 voltage divider, so the gain seen between TP1 and TP2 has this factor folded in).
    • Moreover, the relative scaling of the gain for various slider values also doesn't appear to be liner.
    • For the highest gain setting of +15 dB, the servo began oscillating, so I think the apparent non-flatness of the gain as a function of frequency is an artefact of the measurement.
    • Nevertheless, my conclusion is that the IC should be changed.

I will pull the board and effect the change later today.

I pulled the board out at 345pm after dialling down all the HV supplies in 1X1. I will reinstall it after running some tests.

Attachment 1: VGAchar.pdf  110 kB  | Show | Show all
ELOG V3.1.3-